# Improving Sequential Market Clearing via Value-oriented Renewable Energy Forecasting

### v1

@@ IV-B Loss Function Derivation
In this subsection, we derive the loss function in the neighborhood of the forecast $\hat{\bm{y}}_{d}$. By substituting the functions (21),(23),(24) and dual solutions into the upper-level objective (15a), the loss function in the neighborhood of the forecast $\hat{\bm{y}}_{d}$ is,
@@
Since functions in (26b) are linear, and the dual decision policies are constant, the loss function $\ell_{d}(\hat{\bm{y}}_{d})$ in the neighborhood of the forecast $\hat{\bm{y}}_{d}$ is linear, and output the overall operation cost (4a) given the forecast $\hat{\bm{y}}_{d}$. Naturally, the derivative of $\ell_{d}(\hat{\bm{y}}_{d})$ w.r.t. the forecast $\hat{\bm{y}}_{d}$, i.e., $\frac{\ell_{d}(\hat{\bm{y}}_{d})}{\hat{\bm{y}}_{d}}$, measures the marginal impact of the forecast on the overall cost. It is derived as,
$\begin{split}&(\frac{\ell_{d}(\hat{\bm{y}}_{d})}{\hat{\bm{y}}_{d}})^{\top}=-\bm{\sigma}_{d}^{*\top}\bm{F}_{\text{DA}}^{y}-\bm{\nu}_{d,1}^{*\top}\bm{F}_{\text{RT}}\frac{\partial f_{d,1}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}+\sum_{\tau=2}^{T}-\bm{\zeta}_{d,\tau}^{*T}\\ &(\bm{F}_{\text{RT}}^{\prime x}\frac{\partial f_{d,\tau}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}+\bm{F}_{\text{RT}}^{\prime p}\frac{\partial f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}+\bm{F}_{\text{RT}}^{\prime+-}\frac{\partial f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}})\end{split}$ (27)
According to Theorem 1, the loss function defined across the entire space of $\hat{\bm{y}}_{d}$ is expected to be a piecewise linear function. Specifically, each piece is related with a different active constraint index sets $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau},\forall\tau=1,...,T$ of DA and RT market clearings. Such index sets then determine the primal decision policies (18)-(20), and the following functions in (21)-(25). It is possible to enumerate all possible active constraint index sets, and derive the corresponding loss function [22]. However, implementing such a practice can be computationally expensive, particularly when dealing with large-scale optimization problems. We notice that the derivatives $\frac{\partial f_{d,\tau}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$ in (27) associated with the active index sets are constants, as (21),(23),(24) are linear functions. This suggests that there is no need to recalculate these derivatives when encountering the same active index sets during the training. For that, we propose a solution strategy, where the derivatives are recalculated only when encountering new ones.
@@ V Solution Strategy
We illustrate the training phase of the forecast model based on neural networks (NNs). With the loss function, we use batch optimization to train NN. Given a batch of data in $B$ days, the parameter estimation with the derived loss function is formulated as,
@@
Different from the conventional unconstrained program at the training phase, (28) is with the box constraint (28c) for the NN output. We design a specific model structure to address this. Specifically, a Sigmoid function, whose output is between 0 and 1, is used as the activation function at the output layer. By multiplying its output with the cap $\bar{\bm{y}}_{d,\tau}$ of each sample, the box constraint (28c) is satisfied. NN’s structure is illustrated in Fig. 4.
@@
During the training, the NN, with the parameter $\Theta$ given by any value, outputs the forecast $\hat{\bm{y}}_{d,\tau}$ by (13). The primal and dual problems of DA and RT clearing in (10),(11),(12) are solved. The active constraint index sets $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau},\forall\tau=1,...,T$ and the dual solutions $\bm{\sigma}_{d}^{*},\bm{\nu}_{d,1}^{*},\bm{\zeta}_{d,\tau}^{*},\forall\tau=2,...,T$ are obtained. We check whether they are the new ones. If yes, we calculate derivatives $\frac{\partial f_{d,\tau}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$ associated with the new ones, and calculate the gradient $\frac{\ell_{d}(\hat{\bm{y}}_{d})}{\hat{\bm{y}}_{d}}$ in (27). Additionally, we store the new active index sets and the associated derivatives in the buffers. If not, the stored elements in the buffers are used for calculating the gradient. The training process is summarized in Fig. 5.
Figure 5: The flowchart of the algorithm for the training phase. Require: Learning rate $\alpha$, batch size $B$, initialized buffers $\Lambda_{\text{DA}},\Lambda_{\text{RT},\tau},\forall\tau=1,...,T$ for storing active constraint index set/derivative, and initialized forecast model parameters $\Theta^{(0)}$.
@@ VI Case Study
Table I: Cost and technical data in DA and RT market.
@@
Marginal up-regulation cost $\bm{\rho}_{+}$ (\$/MW) 50 52 54
Marginal down-regulation cost $\bm{\rho}_{+}$ (\$/MW) 18 16 14
@@
We use a four-layer ResNet as the forecast model, which has 256 hidden layer units. Its structure is described in Figure 4. The input context consists of the numeric weather prediction (i.e., the predicted wind speed and direction at 10m and 100m altitude) of each wind farm in the system. We use Root Mean Squared Error (RMSE) on the test set for assessing the forecast quality, and the average overall operation cost on the test set, as defined in (4a), for evaluating the operation value. Four benchmark models are used for comparison: Two quality-oriented, one value-oriented forecasting approach, and a stochastic program. The two quality-oriented forecast models are trained using Mean Squared Error (MSE) and pinball loss (an asymmetric loss function), denoted as Qua-E and Qua-Q, respectively. Specifically, Qua-E provides predictions for expected wind power, while Qua-Q offers quantile predictions. Light Gradient Boosted Machine is used for issuing quantiles, which is the winner of GEFCom 2014 [25]. We consider the value-oriented forecast model trained via OptNet [15] as a benchmark, referred to as OptNet. Lastly, the stochastic program, with 50 wind power scenarios obtained by k-nearest-neighbors, is considered and denoted as Sto-OPT. For each sample on test set, the stochastic program is solved for settling the schedule of generators and wind power in DA. Then, the adjustment in (11) and (12) are performed in RT.
@@ VI-A The Operational Advantage
Since Sto-OPT does not need training or rely on a point forecast, its RMSE is not reported. Sto-OPT serves as the ideal benchmark [9], which has the least average operation cost. The proposed approach outperforms all other methods in terms of average operation cost on the test set, except for Sto-OPT. However, its test time is much shorter than Sto-OPT, demonstrating computational efficiency. Also, we observe that the performance of Sto-OPT is heavily influenced by the number of scenarios used. When fewer scenarios, such as 20, are employed, the average operation cost on the test set increases to \$84,478, which is even worse than that achieved by the proposed approach.
@@
Table II: RMSE, average operation cost on the test set and the training/test time of the proposed approach and benchmarks.
Proposed Qua-E Qua-Q OptNet Sto-OPT
@@ VI-B The Sensitivity Analysis
In this section, we compare the proposed approach against Qua-E under different wind power capacities and adjustment cost $\overline{\bm{\rho}^{+}}$ for up-regulation.
@@ VI-B1 Performance under Different Wind Power Penetration
Figure 7: Average operation cost under different values of wind power capacity. Qua-E stands for quality-oriented forecast issuing expectation.
@@ VI-B2 Performance under Different Up-regulation Cost in RT
To sum up, the operation advantage of the proposed approach is more evident, under large penetration of wind power, and high up-regulation cost.
Table III: Average operation cost in different settings of up-regulation costs in RT. The first column gives the RT up-regulation costs and marginal opportunity loss in each setting. Inside the parenthesis, the up-regulation costs are given on the left of the slash, and the marginal opportunity loss is given on the right.
@@
High up- regulation cost ($G_{1}$: 80/60, $G_{2}$: 82/60, $G_{3}$: 84/60) 85114 92486 8%
Low up- regulation cost ($G_{1}$: 21/1, $G_{2}$: 23/1, $G_{3}$: 25/1) 81517 81677 0.2%
@@ VII Conclusions
We propose a value-oriented renewable energy forecasting approach, for minimizing the expected overall operation cost in the existing deterministic market clearing framework. We analytically derive the loss function for value-oriented renewable energy forecasting in sequential market clearing. The loss function is proved to be piecewise linear when the market clearing is modeled by linear programs. Additionally, we provide the analytical gradient of the loss function with respect to the forecast, which leads to an efficient training strategy. In the case study, compared to quality-oriented forecasting approach trained by MSE, the proposed approach can reduce average operation cost on the test set to 2.9%. Such an advantage is more obvious under large wind power capacity and high up-regulation costs. Under high up-regulation costs, our approach can reduce the cost by up to 8%. Future work will include attempts to derive the loss function for market clearing modeled by other types of optimization programs, such as quadratic optimization or conic optimization.
Appendix A Details of Coefficients in Day-Ahead and Real-Time Clearings
[…]
$\bm{D}_{1}=\begin{bmatrix}\bm{1}^{\top}&\bm{0}^{\top}&\dots&\bm{0}^{\top}\\ \bm{0}^{\top}&\bm{1}^{\top}&\dots&\bm{0}^{\top}\\ \vdots&\vdots&\ddots\\ \bm{0}^{\top}&\bm{0}^{\top}&\dots&\bm{1}^{\top}\end{bmatrix}\bm{D}_{2}=\begin{bmatrix}\tilde{\bm{H}}&\bm{O}&\dots&\bm{O}\\ \bm{O}&\tilde{\bm{H}}&\dots&\bm{O}\\ \vdots&\vdots&\ddots\\ \bm{O}&\bm{O}&\dots&\tilde{\bm{H}}\end{bmatrix}$
$\bm{D}_{3}=\begin{bmatrix}\tilde{\bm{E}}&\bm{O}&\dots&\bm{O}\\ \bm{O}&\tilde{\bm{E}}&\dots&\bm{O}\\ \vdots&\vdots&\ddots\\ \bm{O}&\bm{O}&\dots&\tilde{\bm{E}}\end{bmatrix}\bm{D}_{4}=\begin{bmatrix}-\tilde{\bm{E}}&\tilde{\bm{E}}&\dots&\bm{O}\\ \bm{O}&-\tilde{\bm{E}}&\dots&\bm{O}\\ \vdots&\vdots&\ddots\\ \bm{O}&\bm{O}&\dots&\tilde{\bm{E}}\end{bmatrix}$
$\bm{D}_{5}=\begin{bmatrix}\hat{\bm{E}}&\bm{O}&\dots&\bm{O}\\ \bm{O}&\hat{\bm{E}}&\dots&\bm{O}\\ \vdots&\vdots&\ddots\\ \bm{O}&\bm{O}&\dots&\hat{\bm{E}}\end{bmatrix}\bm{d}_{1}=\begin{bmatrix}\bm{1}^{\top}\bm{l}_{d,1}\\ \vdots\\ \bm{1}^{\top}\bm{l}_{d,T}\end{bmatrix}\bm{d}_{2}=\begin{bmatrix}\bar{\bm{f}}+\bm{H}\bm{l}_{d,1}\\ \vdots\\ \bar{\bm{f}}+\bm{H}\bm{l}_{d,T}\end{bmatrix}$
$\bm{d}_{3}=\begin{bmatrix}\bar{\bm{f}}-\bm{H}\bm{l}_{d,1}\\ \vdots\\ \bar{\bm{f}}-\bm{H}\bm{l}_{d,T}\end{bmatrix}\bm{d}_{4}=\begin{bmatrix}\bar{\bm{p}}\\ \vdots\\ \bar{\bm{p}}\end{bmatrix}\bm{d}_{5}=\begin{bmatrix}-\underline{\bm{p}}\\ \vdots\\ -\underline{\bm{p}}\end{bmatrix}\bm{d}_{6}=\begin{bmatrix}\bar{\bm{r}}\\ \vdots\\ \bar{\bm{r}}\end{bmatrix}\bm{d}_{7}=\begin{bmatrix}-\underline{\bm{r}}\\ \vdots\\ -\underline{\bm{r}}\end{bmatrix}$
$\bm{G}_{\text{DA}}$ and $\bm{\psi}_{\text{DA},d}$ in (10) are the vertical stack of the matrices $\bm{D}_{1}$-$\bm{D}_{5}$ and the vectors of $\bm{d}_{1}$-$\bm{d}_{7}$, i.e., $\bm{G}_{\text{DA}}=[\bm{D}_{1};-\bm{D}_{1};\bm{D}_{2};-\bm{D}_{2};\bm{D}_{3};-\bm{D}_{3};\bm{D}_{4};-\bm{D}_{4};\bm{D}_{5};-\bm{D}_{5}]$ and $\bm{\psi}_{\text{DA},d}=[\bm{d}_{1};-\bm{d}_{1};\bm{d}_{2};\bm{d}_{3};\bm{d}_{4};\bm{d}_{5};\bm{d}_{6};\bm{d}_{7};\bm{0};\bm{0}]$. The matrix $\bm{F}_{\text{DA}}^{y}$ is the vertical stack of all-zero matrix and identity matrix, i.e., $\bm{F}_{\text{DA}}^{y}=[\bm{O};...;\bm{O};\bm{E};\bm{O}]$. Let $\tilde{\bm{\rho}}=[\bm{\rho};\bm{0}]$. Then, $\bm{\rho}_{\text{DA}}=[\tilde{\bm{\rho}};...;\tilde{\bm{\rho}}]$.
Next, the details of coefficients in RT clearing (11) are provided. We define the following matrices, which are used for forming the coefficients in (11),
$\bm{R}_{1}=\begin{bmatrix}\bm{1}^{\top}&-\bm{1}^{\top}&-\bm{1}^{\top}\end{bmatrix},\bm{R}_{2}=\begin{bmatrix}\bm{H}&-\bm{H}&-\bm{H}\end{bmatrix}$
$\bm{R}_{3}=\begin{bmatrix}\bm{E}&\bm{O}&\bm{O}\end{bmatrix},\bm{R}_{4}=\begin{bmatrix}\bm{O}&\bm{E}&\bm{O}\end{bmatrix}$
$\bm{R}_{5}=\begin{bmatrix}\bm{O}&\bm{O}&\bm{E}\end{bmatrix},\bm{r}_{2}=\bar{\bm{f}}+\bm{H}(\bm{l}_{d,\tau}-\bm{y}_{d,\tau})$
$\bm{r}_{3}=\bar{\bm{f}}+\bm{H}(-\bm{l}_{d,\tau}+\bm{y}_{d,\tau}),\bm{r}_{4}=\overline{\bm{p}^{+}}$
$\bm{r}_{5}=\underline{\bm{p}^{-}},\bm{r}_{6}=\overline{\bm{p}}-\bm{p}_{d,\tau}^{*},\bm{r}_{7}=\bm{p}_{d,\tau}^{*}-\underline{\bm{p}},\bm{r}_{8}=\bm{y}_{d,\tau}$
$\bm{G}_{\text{RT}}$ and $\bm{\psi}_{\text{RT},d,\tau}$ in (11) are the vertical stack of the matrices of $\bm{R}_{1}$-$\bm{R}_{5}$ and the vectors of $\bm{r}_{1}$-$\bm{r}_{8}$. $\bm{G}_{\text{RT}}=[\bm{R}_{1};-\bm{R}_{1};\bm{R}_{2};-\bm{R}_{2};\bm{R}_{3};-\bm{R}_{3};\bm{R}_{4};-\bm{R}_{4};\bm{R}_{3};-\bm{R}_{3};$ $\bm{R}_{4};-\bm{R}_{4};\bm{R}_{5};-\bm{R}_{5}]$ and $\bm{\psi}_{\text{RT},d,\tau}=[r_{1};-r_{1};\bm{r}_{2};\bm{r}_{3};\bm{r}_{4};\bm{0};$ $\bm{r}_{5};\bm{0};\bm{r}_{6};\bm{0};\bm{r}_{7};\bm{0};\bm{r}_{8};\bm{0}]$. Let $\bm{I}_{\text{RT}}=[\bm{0},\bm{1}^{\top}]$ and $\bm{H}_{\text{RT}}=[\bm{H},\bm{O}]$. Then, $\bm{F}_{\text{RT}}=[\bm{I}_{\text{RT}};-\bm{I}_{\text{RT}};\bm{H}_{\text{RT}};-\bm{H}_{\text{RT}};\bm{O};...;\bm{O};-\tilde{\bm{E}};\bm{O};\tilde{\bm{E}};\bm{O};\bm{O};\bm{O}]$. $\bm{\rho}_{\text{RT}}=[\bm{\rho}_{+};-\bm{\rho}_{-}]$.
Finally, the details of coefficients in RT clearing (12) are provided. $\bm{G}^{\prime}_{\text{RT}}=[\bm{R}_{1};-\bm{R}_{1};\bm{R}_{2};-\bm{R}_{2};\bm{R}_{3};-\bm{R}_{3};\bm{R}_{4};-\bm{R}_{4};$ $\bm{R}_{3};-\bm{R}_{3};\bm{R}_{4};-\bm{R}_{4};\bm{R}_{3};-\bm{R}_{3};\bm{R}_{4};-\bm{R}_{4};\bm{R}_{5};-\bm{R}_{5}]$ and $\bm{\psi}^{\prime}_{\text{RT},d,\tau}=[r_{1};-r_{1};\bm{r}_{2};\bm{r}_{3};\bm{r}_{4};\bm{0};\bm{r}_{5};\bm{0};\bm{r}_{6};\bm{0};\bm{r}_{7};\bm{0};\bar{\bm{r}};\bm{0};-\bar{\bm{r}};\bm{0};$ $\bm{r}_{8};\bm{0}]$. Then, $\bm{F}_{\text{RT}}^{\prime x}=[\bm{I}_{\text{RT}};-\bm{I}_{\text{RT}};\bm{H}_{\text{RT}};-\bm{H}_{\text{RT}};\bm{O};...;\bm{O};-\tilde{\bm{E}};\bm{O};$ $\tilde{\bm{E}};\bm{O};-\tilde{\bm{E}};\bm{O};\bm{O};\bm{O}]$. $\bm{F}_{\text{RT}}^{\prime p}=[\bm{O};...;\bm{E};\bm{O};-\bm{E};\bm{O};\bm{O};\bm{O}]$. Let $\bm{E}_{\text{RT}}=[\bm{E},-\bm{E}]$. Then, $\bm{F}_{\text{RT}}^{\prime+-}=[\bm{O};...;\bm{E}_{\text{RT}};\bm{O};-\bm{E}_{\text{RT}};\bm{O};\bm{O};\bm{O}]$.

### v2

@@ IV-B Loss Function Derivation
By substituting the functions (21),(23),(24) and dual solutions into the upper-level objective (15a), the loss function in the neighborhood of the forecast $\hat{\bm{y}}_{d}$ is,
@@
Since functions in (26b) are linear, and the dual decision policies are constant, the loss function $\ell_{d}(\hat{\bm{y}}_{d})$ in the neighborhood of the forecast $\hat{\bm{y}}_{d}$ is linear, and outputs the overall operating cost (4a) given the forecast $\hat{\bm{y}}_{d}$. Naturally, the derivative of $\ell_{d}(\hat{\bm{y}}_{d})$ w.r.t. the forecast $\hat{\bm{y}}_{d}$, i.e., $\frac{\partial\ell_{d}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, measures the marginal impact of the forecast on the overall cost and is a constant. It is,
$\begin{split}&(\frac{\partial\ell_{d}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}})^{\top}=-\bm{\sigma}_{d}^{*\top}\bm{F}_{\text{DA}}^{y}-\bm{\nu}_{d,1}^{*\top}\bm{F}_{\text{RT}}\frac{\partial f_{d,1}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}+\sum_{\tau=2}^{T}-\bm{\zeta}_{d,\tau}^{*T}\\ &(\bm{F}_{\text{RT}}^{\prime x}\frac{\partial f_{d,\tau}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}+\bm{F}_{\text{RT}}^{\prime p}\frac{\partial f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}+\bm{F}_{\text{RT}}^{\prime+-}\frac{\partial f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}})\end{split}$ (27)
where the derivatives $\frac{\partial f_{d,\tau}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$ in (27) are constants, which measure the marginal impact of the forecast on the primal solutions. We have the following proposition regarding the loss function.
Proposition 3.
The loss function defined across the entire space of $\hat{\bm{y}}_{d}$ is a piecewise linear function. Each piece is a linear function described in (26).
Proof.
According to Theorem 1, (21),(23),(24) are piecewise linear functions defined across the entire space of $\hat{\bm{y}}_{d}$. Since the loss function is the linear transformation of (21),(23),(24), it is also a piecewise linear function. ∎
Since (21),(23),(24) are determined by different active constraint index sets $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau-1},\forall\tau=2,...,T$ of DA and RT market clearing, each piece of the loss function is related with different active constraint index sets we well. One way is to enumerate all possible active constraint index sets, and derive the corresponding piece of the loss function [23]. However, the enumeration can be computationally expensive, particularly when dealing with large-scale optimization.
Therefore, we propose to derive the pieces of the loss function in an on-demand manner. Concretely, during training, we only recalculate the piece of the loss function when encountering new active index sets. The details are given in the next Section.
@@ V Solution Strategy
We illustrate the training phase of the forecasting model based on neural networks (NNs). With the loss function, we use batch optimization to train NN. Given a batch of data in $B$ days, the parameter estimation with the derived loss function is formulated as,
@@
Different from the conventional unconstrained program at the training phase, (28) is with the box constraint (28c) for the NN output. We design a specific model structure to address this. Specifically, a Sigmoid function, whose output is between 0 and 1, is used as the activation function at the output layer. By multiplying its output with the cap $\bar{\bm{y}}_{d,\tau}$ of each sample, the constraint (28c) is satisfied. NN’s structure is in Fig. 5.
@@
During the training, the NN, with the parameter $\Theta$ given by any value, outputs the forecast $\hat{\bm{y}}_{d,\tau}$ by (13). The primal and dual problems of DA and RT clearing in (10),(11),(12) are solved. The active constraint index sets $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau-1},\forall\tau=2,...,T$ and the dual solutions $\bm{\sigma}_{d}^{*},\bm{\nu}_{d,1}^{*},\bm{\zeta}_{d,\tau}^{*},\forall\tau=2,...,T$ are obtained. We check if the active constraint index sets are in the buffers $\Lambda_{\text{DA}}\Lambda_{\text{RT},\tau-1},\forall\tau=2,...,T$. If yes, the stored derivatives $\frac{\partial f_{d,\tau}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$ in the buffers are used for calculating the gradient $\frac{\partial\ell_{d}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$ in (27). If not, we calculate the derivatives associated with the new ones, and calculate the gradient in (27). Additionally, we store the new active index sets and the associated derivatives in the buffers. The illustration of gradient calculation is in Fig. 6.
Figure 6: The illustration of gradient calculation in the training process.
@@ VI Case Study
Table I: Cost and technical data in DA and RT market of IEEE 9-Bus system.
@@
Marginal cost $\bm{\rho}_{+}$ (\$/MW) 50 52 54
Marginal utility $\bm{\rho}_{-}$ (\$/MW) 18 16 14
@@
We use a four-layer ResNet as the forecasting model, which has 256 hidden layer units. Its structure is described in Figure 5. The input context consists of the numeric weather prediction (i.e., the predicted wind speed and direction at 10m and 100m altitude) of each wind farm in the system. We use Root Mean Squared Error (RMSE) on the test set for assessing the forecast quality, and the average overall operating cost on the test set, as defined in (4a), for evaluating the operation value. Four benchmark models are used for comparison: Two quality-oriented (Qua-E and Qua-Q), one value-oriented point forecasting approach (OptNet), and a stochastic market clearing (Sto-OPT-P), i.e.,
  1. 1.
    Qua-E: The forecasting model (ResNet) is trained with the loss function of Mean Squared Error (MSE). The trained model predicts conditional expectations.
  2. 2.
    Qua-Q: The forecasting model is trained with the pinball loss. The trained model offers quantile predictions. Light Gradient Boosted Machine, the winner of GEFCom 2014 [25], is used as the forecasting model.
  3. 3.
    OptNet: The forecasting model (ResNet) is trained by differentiable programming [15, 14], to minimize overall system operating costs.
  4. 4.
    Sto-OPT-P: For stochastic clearing, 50 wind power scenarios are obtained via k-nearest-neighbors, which is a Prescriptive approach [26]. For each sample on the test set, the stochastic clearing is for settling the schedule of generators and wind power in DA. Then, the adjustment in (11) and (12) are performed in RT.
@@ VI-A The Operational Advantage
Since Sto-OPT-P does not need training or relies on a point forecast, its RMSE is not reported. Sto-OPT-P serves as the ideal benchmark [7], which has the least average operating cost. The proposed approach outperforms all other methods in terms of average operating cost on the test set, except for Sto-OPT-P. However, its test time is much shorter than Sto-OPT-P, demonstrating computational efficiency. Also, we observe that the performance of Sto-OPT-P is heavily influenced by the number of scenarios used. When fewer scenarios, such as 20, are employed, the average operating cost on the test set increases to \$84,478, which is even worse than that achieved by the proposed approach.
@@
Table II: RMSE, average operating cost on the test set and the training/test time of the proposed approach and benchmarks.
Proposed Qua-E Qua-Q OptNet Sto-OPT-P
@@ VI-B The Sensitivity Analysis
In this section, we compare the proposed approach against Qua-E under different wind power capacities and marginal cost $\bm{\rho}^{+}$ for up-regulation.
@@ VI-B1 Performance under Different Wind Power Penetration
Figure 8: Average operating cost under different wind power capacities. Qua-E stands for quality-oriented forecast issuing expectation.
@@ VI-B2 Performance under Different Up-regulation Cost in RT
To sum up, the operation advantage of the proposed approach is more evident, under large penetration of wind power, and high up-regulation marginal cost.
Table III: Average operating cost in different settings of up-regulation costs in RT. The first column gives the RT up-regulation costs and marginal opportunity loss in each setting. Inside the parenthesis, the up-regulation marginal costs are given on the left of the slash, and the marginal opportunity loss is given on the right.
@@
High up- regulation marginal cost ($G_{1}$: 80/60, $G_{2}$: 82/60, $G_{3}$: 84/60) 85114 92486 8%
Low up- regulation marginal cost ($G_{1}$: 21/1, $G_{2}$: 23/1, $G_{3}$: 25/1) 81517 81677 0.2%
@@
VI-C Extending to Larger Scale System
We test the proposed approach on the IEEE 39-bus system, with 21 loads, 2 wind farms at buses 38 and 39, and 10 generators. The illustration of the IEEE 39-bus system and parameters for generators are given in our Github repo [24]. The average operating costs of the proposed one and Qua-E are shown in Fig. 8(b). The proposed approach results in lower average operating costs than Qua-E. With increased wind power capacity in the IEEE 39-bus system, the advantage of the proposed approach in decreasing the overall costs is even more pronounced compared to the IEEE 9-bus system.
@@ VII Conclusions
We propose a value-oriented renewable energy forecasting approach, for minimizing the expected overall operating cost in the existing deterministic market clearing framework. We analytically derive the loss function for value-oriented renewable energy forecasting in sequential market clearing. The loss function is proved to be piecewise linear when the market clearing is modeled by linear programs. Additionally, we provide the analytical gradient of the loss function with respect to the forecast, which leads to an efficient training strategy. In the case study, compared to quality-oriented forecasting approach trained by MSE, the proposed approach can reduce average operating cost on the test set to 2.9% for the IEEE 9-bus system. Such an advantage is more obvious under large wind power capacity and high up-regulation costs. Under high up-regulation marginal costs, our approach can reduce the cost by up to 8%. Future work will include derivation of the value-oriented loss function for market clearing modeled by other types of optimization programs, such as quadratic optimization and conic optimization.